solve. 4(7x+10)=3(9x-5)+20
Answer by checkley71(8403)

4(7X+10)=3(9X-5)+20
28X+40=27X-15+20 COMBINE ALL X TERMS ON THE LEFT SIDE & ALL NUMBERS ON THE RIGHT SIDE
28X-27X=-15+20-40
X=-35 SOLUTION
PROOF
4(7*-35+10)=3(9*-35-5)+20
4(-245+10)=3(-315-5)+20
4(-235)=3(-320)+20
-940=-960+20
-940=-940